OpenStudy (anonymous):

9g contains ________ molecules..??

OpenStudy (anonymous):

9 g of what?!

OpenStudy (anonymous):

9g of water

OpenStudy (anonymous):

n= m/M n=N/L so N= n*L N= m/m * L

OpenStudy (anonymous):

i made mistake N= (m/M) * L

OpenStudy (anonymous):

what value?

OpenStudy (anonymous):

methametical value..??

OpenStudy (anonymous):

3,011 * 10^23 molecules

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem